A principal of $3,000 is invested at 5.75% interest compounded annually how much will the investment be worth after 11 years
Answer by rapture1965(18)

You need the formula A = (1 + r)^(t), where the following applies:

A = compounded annually
r = rate
t = time (in years)
p = principal amount

Just plug the information in the problem into the formula and solve for A.

A = 3000(1 + 0.0575)^(11)

A = 3000(1.0575)^(11)

A = 3000(1.8496269153)

A = 5,548.8807457559

We can round this decimal number to the first decimal place.

A = 5,548.90

After 11 years, the investment will be worth $5,548.90.